# Geometric Framework for Cell Oversegmentation

### v2

@@ 3.3 Method Generalization
Recall that our original framework operates on oversegmentation results that the cutting surfaces between oversegmented cell pairs are always aligned with the horizontal plane of the standard coordinate axes. However, in non-2D-based approaches, these cutting surfaces are often tilted, irregular, and rugged. To address this, we extend our framework to effectively handle these more complex oversegmentation scenarios. We validate this generalized method on the Pancreas-B dataset, where the non-2D-based approach demonstrates better performance (see Figure 16), likely because the cells in Pancreas-B resemble those in the ovule dataset used to train PlantSeg. Notably, we observe that tilted oversegmentation is more prevalent in this dataset than the others. We also propose several recommendations to improve the end-user experience in such settings.
@@
Candidates Screening. Since oversegmentation can occur in any direction for 3D-based approaches, we store all cell information in a graph structure and retrieve neighboring cells—i.e., those in direct contact—as suspected candidates. The number of suspected candidates in this setting is expected to be significantly larger than in the axis-aligned case described in §3.2.
@@
Oversegmentation Cutting Surface. We define the cutting surface between two cells as the shared curved interface where the cells are in direct contact. In the previous setting, this surface was aligned with the plane, and 2D segmentation layers were sliced from the 3D volume along this standard plane. In this generalized case, we extend this concept by slicing the 3D segmentation volume layer-by-layer along a rotated reference plane—one that is aligned with the cutting surface between two oversegmented cells—thereby forming a rotated 2D segmentation representation.
@@
Rotated 3D Segmentation Reconstruction. The first step is to estimate a reference plane from the rugged cutting surface between two oversegmented cells. We treat this cutting surface as a curvature and apply PCA to extract a best-fit plane, which serves as the estimated reference plane. Using this reference plane, we slice the 3D segmentation volume into parallel 2D layers by shifting the plane upward and downward. However, as illustrated in Step 4 of Figure 5, since all cells are stored in a discrete index format, the tilted plane intersects each layer as a series of discrete strips (aligned with integer indices), rather than forming smooth, continuous 2D geometries. To address this, we interpolate between these strips to reconstruct complete 2D cell masks. The reconstructed 2D masks are then used as input to our pretrained classifier. Specifically, we remove the masks of both cells at the reference layer (i.e., the unshifted PCA plane) and treat the remaining upward and downward segments as a pair of candidate cell fragments. The classifier then determines whether these two parts should be merged. Notably, in this setting, layer-to-layer interpolation is no longer required to correct 2D mis-segmentation—if merging is validated, we directly combine the original 3D cell bodies.
Practical Scheme. Although 2D-based and 2.5D-based approaches remain the mainstream in practice, our generalized framework aims to support a broader range of use cases for non-2D-based approaches. To provide end-users with greater control and flexibility, we introduce tunable hyperparameters into this generalized framework. Technical details of the implementation are provided in Appendix B.2.
4 Experiment
@@
Dataset Overview. We evaluate our method on two types of datasets: publicly available, labeled plant cell datasets, and private, unlabeled animal cell datasets. Specifically, the animal datasets lack ground-truth annotation. In total, we use six different plant cell datasets, each consisting of 100 image stacks, and two animal cell datasets with 12 image stacks. Notably, each animal image stack is a large volumetric dataset containing approximately 300 image layers, with the maximum plane size reaching up to pixels. After the initial 3D segmentation, each image stack contains approximately 2,000 to 4,000 cells. Further dataset details are provided in Appendix D.
@@
Experiment Overview. Our experimental evaluation consists of three parts: (i) synthesized test cases derived from ground-truth plant datasets, which provide deterministic labels and enable accurate evaluation using metrics such as recall and F1-score; (ii) direct augmentation of real, flawed 3D plant segmentation results, which we evaluate it across multiple state-of-the-art methods using both deterministic quantitative metrics and qualitative results validated by human experts; (iii) direct augmentation of flawed 3D segmentation results on animal datasets, which serves to assess the transferability of our method in the absence of ground truth; and (iv) ablation analysis on the Geo-Wasserstein divergence—specifically, evaluating the impact of removing layer-to-layer EMD information on the final results. We believe this experimental design and evaluation setup provide strong justification for the effectiveness and robustness of our approach.
4.1 Plant Cells: Synthesized Cases
In this task, we reserve a small portion (10 image stacks) from each plant dataset, containing true (oversegmented) and false (natural gap) test cases that are excluded from the training set. This setup allows us to evaluate the pretrained model on unseen but similar data. Since these test cases are constructed from labeled ground-truth segmentations, each has a deterministic ground-truth label, enabling precise quantitative evaluation. Results are shown in Table 2:
| Dataset | Sepal | Pedicel | Valve | Leaf | Anther | Filament |
| Recall | 0.955 | 0.854 | 0.965 | 0.988 | 0.840 | 0.854 |
| Precision | 0.963 | 0.998 | 0.983 | 1.000 | 0.985 | 0.981 |
| F1-Score | 0.959 | 0.920 | 0.974 | 0.994 | 0.907 | 0.913 |
However, the accuracy reported above serves only as a reference for the pretrained model’s performance under ideal conditions with perfect initial pre-segmented 2D results, but in pratice, pre-segmentation inaccuracies are inevitable. Therefore, we next evaluate the actual performance of our method on real, flawed 3D segmentation results generated by CellPose, CellStitch, and PlantSeg.
4.2 Plant Cells: Real Cases
In this task, we apply CellPose, CellStitch, and PlantSeg directly to the raw microscopy images to obtain the initial 3D segmentation results, which inherently contain errors due to imperfect segmented cell masks—such as shrinkage and fragmentation—commonly encountered in real-world applications. We then apply our proposed method to these flawed segmentations. This experiment is designed to evaluate the robustness of our framework in handling realistic segmentation settings.
Evaluation Metrics. We first report the quantitative results to directly assess the improvement in overall segmentation quality using standard evaluation metrics from multiple perspectives: mean Average Precision (mAP) and Jaccard Index (Jac). Specifically, the average precision (AP) is computed as , where , , and represent the numbers of true positive, false negative, and false positive masks, respectively, under a specified Intersection over Union (IoU) threshold . The mean Average Precision (mAP) is then obtained by averaging the AP scores at IoU thresholds of , , and . The Jaccard Index (Jac), by contrast, directly measures the IoU overlap between predicted and ground truth cell masks. Results are shown in Table 3.
Necessary Case-by-case Validation. However, as a new formulated problem, we would like to discuss the limitations and biases of these standard metrics in our task, and why case-by-case human evaluation remains necessary: (i) mAP may decrease even when the correction is correct. In Figure 11 (a), we show a case where two oversegmented masks were correctly stitched, yet the mAP score decreased after correction. This is because mAP and other overlap-precision-based metrics rely on a fixed IoU threshold between predicted and ground truth cell masks. During post-processing, segmentation methods often shrink the original masks, so our corrected mask at layer , built from these already shrunken masks, may fall below the IoU threshold when compared to the ground truth; (ii) Jaccard may increase even when the correction is incorrect. Although Jaccard index shows significant improvement across multiple datasets, we highlight that incorrect corrections can also inflate this metric. Shown in Figure 11 (b), the correction is actually wrong when closely compared to the ground truth masks. This issue affects all other overlapping-area-based metrics as well.
@@ 4.2 Plant Cells: Real Cases
Therefore, while these quantitative metrics are important and do indicate general improvement, we emphasize that human case-by-case evaluation is still essential for a complete understanding of the correction quality. During human verification, we classify oversegmentation corrections into three categories: (i) Correct cases, where the oversegmented cell masks are present in the ground truth and belong to the same cell; (ii) Incorrect cases, where the oversegmented cell masks are present in the ground truth but actually belong to different cells (natural gap); and (iii) Unsure cases, which include ambiguous scenarios such as “hallucination”, where the correction appears geometrically plausible but lacks definitive support from the ground truth. A common source of such cases is the hallucination effect in CellPose2D (see Figure 12 and Figure 13), where highlight spots or noise in the raw image are mistakenly interpreted as valid cell masks. Although these regions are empty in the ground truth, they are treated as real masks by 3D models. Consequently, such artifacts propagate into the 3D segmentation and affect our correction process, leading to visually “correct” results that are based on flawed inputs. An example of the correct case is shown in Figure 6, where in the CellStitch row, layers and contain cell masks (highlighted using the green box) that are also appeared in the ground truth row at the same position. Our method detects these two masks and stitches them as the same cell. We provide more visualization examples of unsure and incorrect cases in Appendix C.3 and Appendix C.4.
Results Discussion. By analyzing the incorrect cases, we identified a major causes of these errors: CellPose2D often “shrinks the cell mask smaller” during 2D segmentation, especially when the cell boundaries in the raw images are blurred by a light band. This shrinkage results in the loss of geometric information, as the EMD between the two masks is also affected when one mask shrinks.
@@ 4.3 Animal Cells: Real Cases
In this task, we evaluate the transfer learning effectiveness of our pre-trained plant cell model on an unlabeled animal cell dataset (Pancreas-A, abbreviated as PA). We recommend users follow the pipeline in §3.2 to train models on their own data, but still, we offers a reference for performance on unseen and distinct cell types. Due to the lack of ground truth, we present the augmented results on CellStitch and report case-by-case verification result. Results are shown in Table 4:
| Dataset | PA-11 | PA-16 | PA-03 | PA-07 | PA-04 | PA-09 | PA-10 | PA-05 | PA-02 | PA-14 | PA-08 | Overall |
| # Correct | 9 | 10 | 19 | 29 | 21 | 25 | 35 | 52 | 28 | 52 | 94 | 374 |
| # Unsure | 1 | 3 | 3 | 5 | 3 | 9 | 8 | 7 | 8 | 21 | 24 | 92 |
| # Incorrect | 1 | 2 | 5 | 4 | 3 | 6 | 4 | 10 | 1 | 2 | 14 | 52 |
| Accuracy | 0.900 | 0.833 | 0.792 | 0.879 | 0.875 | 0.807 | 0.897 | 0.839 | 0.966 | 0.963 | 0.870 | 0.878 |
@@
Results Discussion. Together with the results in §4.2, we observe that the performance of our framework is influenced by the resolution of the image stack and the average size of the cell masks. Higher resolutions amplify geometric details, enabling Geo-Wasserstein divergence to capture more nuanced information. We also include results on the Pancreas-B dataset in Appendix B.2, which exhibits tilted oversegmentation errors when using non-2D-based approaches such as PlantSeg.
4.4 Abation Analysis on Geo-Wasserstein Divergence
| Dataset | Leaf | Sepal | Pedicel | Anther | Filament | Valve | ||||||
| w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | |
| # Correct | 52 | 42 | 126 | 70 | 78 | 71 | 113 | 93 | 139 | 92 | 91 | 75 |
| # Unsure | 64 | 38 | 312 | 113 | 477 | 134 | 371 | 141 | 338 | 124 | 148 | 82 |
| # Incorrect | 36 | 3 | 131 | 13 | 128 | 15 | 102 | 13 | 99 | 28 | 40 | 4 |
| Accuracy | 0.5909 | 0.9333 | 0.4903 | 0.8434 | 0.3786 | 0.8256 | 0.5256 | 0.8807 | 0.5840 | 0.7667 | 0.6945 | 0.9494 |
@@ 4.4 Abation Analysis on Geo-Wasserstein Divergence
We trained the classifier with and without EMD features and evaluated both on CellStitch’s flawed 3D segmentation results for labeled plant cells. Case-by-case verification results are reported in Table 5. Incorporating EMD significantly improved correction accuracy by capturing geometric differences between 2D masks that area-based metrics often overlook. While 3D topological features, such as layer-to-layer overlaps, offer limited information, geometry-aware EMD reliably detects subtle layer-to-layer 2D mask shape variations critical for distinguishing true oversegmentations from natural gaps. For example, two geometries may have identical areas but entirely different shapes, then EMD could capture this shape divergence, while the area-based metrics failed. By combining 2D geometric features with 3D topological information, our framework more effectively captures geometric-topological consistency to identify oversegmentation.
@@ 5 Conclusion
In this work, we present a geometric framework to address 3D cell oversegmentation, an important yet unsolved challenge that severely affects the quality of 3D segmentation results. Our method is designed to identify and correct oversegmented regions by leveraging both 2D geometric and 3D topological information, enabling more accurate reconstruction of cellular structures. We conduct extensive experiments on both synthesized and real datasets to validate the effectiveness of our approach across in-domain and out-of-domain settings. We believe this work has meaningful implications in two areas. First, our proposed framework introduces a new direction for future research in 3D segmentation, particularly using geometry-aware methods to handle complex cellular morphologies, such as neurons and cancer cells with chaotic spatial patterns. Second, we introduce a novel metric, the Geo-Wasserstein divergence, that may inspire future work in geometry processing and low-dimensional geometric representation learning. We provide further discussion of related work in Appendix A, technical details and computational efficiency of the pipeline in Appendix B, and additional visualization results with case-specific analyses in Appendix C.
@@ Acknowledgement
PC and YL would also like to express our appreciation to Andrew Blumberg, Abigail Hickok, Yinuo Jin for their valuable feedback and discussions.

### v3

@@ 3.3 Method Generalization
Recall that our original framework operates on oversegmentation results where the cutting surfaces between cell pairs are always aligned with the horizontal planes of the standard coordinate axes. In contrast, non-2D-based approaches often produce cutting surfaces that are tilted, irregular, or rugged. To address this discrepancy, we extend our framework to accommodate these more complex oversegmentation cases, illustrated in Figure 5.
Candidates Screening.
@@
Since oversegmentation can occur in any direction for 3D-based approaches, we store all cell information in a graph structure and retrieve neighboring cells (i.e., those in direct contact) as suspected candidates. The number of suspected candidates in this setting is expected to be significantly larger than in the axis-aligned case described in §3.2.
Oversegmentation Cutting Surface.
@@
We define the cutting surface between two cells as the shared curved interface where the cells are in direct contact. In the previous setting, this surface was aligned with the plane, and 2D segmentation layers were sliced from the 3D volume along this standard plane. In this generalized case, we extend this concept by slicing the 3D segmentation volume layer-by-layer along a rotated reference plane (one that is aligned with the cutting surface between two oversegmented cells) thereby forming a rotated 2D segmentation representation.
Rotated 3D Segmentation Reconstruction.
@@
The first step is to estimate a reference plane from the rugged cutting surface between two oversegmented cells. We treat this cutting surface as a curvature and apply PCA to extract a best-fit plane, which serves as the estimated reference plane. Using this reference plane, we slice the 3D segmentation volume into parallel 2D layers by shifting the plane upward and downward. However, as illustrated in Step 4 of Figure 5, since all cells are stored in a discrete index format, the tilted plane intersects each layer as a series of discrete strips (aligned with integer indices), rather than forming smooth, continuous 2D geometries. To address this, we interpolate between these strips to reconstruct complete 2D cell masks. The reconstructed 2D masks are then used as input to our pretrained classifier. Specifically, we remove the masks of both cells at the reference layer (i.e., the unshifted PCA plane) and treat the remaining upward and downward segments as a pair of candidate cell fragments.
The classifier then determines whether these two parts should be merged. Notably, in this setting, layer-to-layer interpolation is no longer required to correct 2D mis-segmentation—if merging is validated, we directly combine the original 3D cell bodies. We evaluate these cases specifically for PlantSeg. Results are shown in Table 2 (a).
4 Experiments
Dataset Overview.
@@ 4 Experiments
We evaluate our method on two types of datasets: publicly available, labeled plant cell datasets, and private, unlabeled animal cell datasets. Specifically, the animal datasets lack ground-truth annotation. In total, we use six plant cell datasets [2], each consisting of 100 image stacks, and an animal cell datasets with 11 image stacks. Notably, each animal image stack is a large volumetric dataset containing approximately 300 image layers, with the maximum plane size reaching up to pixels. After the initial 3D segmentation, each image stack contains approximately 2,000 to 4,000 cells. Note that due to the scarcity of high-quality and labeled 3D volumetric image stacks, plant datasets from [2] are commonly used as a standard benchmark for 3D cell segmentation training, as adopted by state-of-the-art methods [17, 14, 35, 31]. To demonstrate the effectiveness under transfer learning, we additionally evaluated on our private animal cell datasets without re-training, showing its applicability across different cell types in organ-level image stacks. Further dataset details are provided in Appendix D.
Experiment Overview.
@@
Our experimental evaluation consists of four parts: (i) synthesized test cases derived from ground-truth plant datasets, which provide deterministic labels and enable accurate evaluation using metrics such as recall and F1-score; (ii) direct augmentation of real, flawed 3D plant segmentation results, which we evaluate it across multiple state-of-the-art methods using both deterministic quantitative metrics and qualitative results validated by human experts; (iii) direct augmentation of flawed 3D segmentation results on animal datasets, which serves to assess the transferability of our method in the absence of ground truth; and (iv) ablation analysis on the key design: the robustness of the statistical feature extraction and the contribution brought by the Geo-Wasserstein divergence.
@@
| Dataset | Sepal | Pedicel | Valve | Leaf | Anther | Filament |
| Recall | 0.955 | 0.854 | 0.965 | 0.988 | 0.840 | 0.854 |
| Precision | 0.963 | 0.998 | 0.983 | 1.000 | 0.985 | 0.981 |
| F1-Score | 0.959 | 0.920 | 0.974 | 0.994 | 0.907 | 0.913 |
| Dataset | PA-11 | PA-16 | PA-03 | PA-07 | PA-04 | PA-09 |
| # Correct | 9 | 10 | 19 | 29 | 21 | 25 |
| # Unsure | 1 | 3 | 3 | 5 | 3 | 9 |
| # Incorrect | 1 | 2 | 5 | 4 | 3 | 6 |
| Accuracy | 0.900 | 0.833 | 0.792 | 0.879 | 0.875 | 0.807 |
| Dataset | PA-10 | PA-05 | PA-02 | PA-14 | PA-08 | Overall |
| # Correct | 35 | 52 | 28 | 52 | 94 | 374 |
| # Unsure | 8 | 7 | 8 | 21 | 24 | 92 |
| # Incorrect | 4 | 10 | 1 | 2 | 14 | 52 |
| Accuracy | 0.897 | 0.839 | 0.966 | 0.963 | 0.870 | 0.878 |
4.1 Plant Cells: Synthesized Cases
In this task, we reserve a small portion (10 stacks) from each plant dataset, containing true (oversegmented) and false (natural gap) test cases that are excluded from the training set. This setup allows us to evaluate the pretrained model on unseen but similar data. Since these test cases are constructed from labeled ground-truth segmentations, each has a deterministic ground-truth label, enabling precise quantitative evaluation. Results are shown in Table 2 (b).
4.2 Plant Cells: Real Cases
In this task, we apply CellPose, CellStitch, and PlantSeg directly to the raw microscopy images to obtain the initial 3D segmentation results, which inherently contain errors due to imperfect segmented cell masks, such as shrinkage and fragmentation, commonly encountered in real-world applications. We then apply our method across §3.2 and §3.3 to these flawed segmentations. This experiment is designed to evaluate the robustness of our framework in realistic segmentation settings with errors from pre-2D results.
Evaluation Metrics.
We first report the quantitative results to directly assess the improvement in overall segmentation quality using standard evaluation metrics from multiple perspectives: mean Average Precision (mAP) and Jaccard Index (Jac). Specifically, the average precision (AP) is computed as , where , , and represent the numbers of true positive, false negative, and false positive masks, respectively, under a specified Intersection over Union (IoU) threshold . The mean Average Precision (mAP) is then obtained by averaging the AP scores at IoU thresholds of . The Jaccard Index (Jac), by contrast, directly measures the IoU overlap between predicted and ground truth cell masks. Results are shown in Table 2 (a).
Apart from quantitative results that reflects the overall segmentation quality improvement, we further incorporate case-by-case validation from human expert to get a deeper understanding of the specific correction quality. During human verification, we classify oversegmentation corrections into three categories: (i) Correct cases, where the oversegmented cell masks are present in the ground truth and belong to the same cell; (ii) Incorrect cases, where the oversegmented cell masks are present in the ground truth but actually belong to different cells (natural gap); and (iii) Unsure cases, which include ambiguous scenarios such as “hallucination”, where the correction appears geometrically plausible but lacks definitive support from the ground truth. We provide further discussion and visualization examples for these cases in Appendix C. An example of the correct case is shown in Figure 6, where in the CellStitch row, layers and contain cell masks (highlighted using the green box) that are also appeared in the ground truth row at the same position. Our method detects these two masks and stitches them as the same cell.
@@ 4.3 Animal Cells: Real Cases
In this task, we evaluate the transfer learning effectiveness of our pre-trained plant cell model on an unlabeled animal pancreas cell dataset. Due to the lack of ground truth, we present the augmented results on CellStitch and report case-by-case verification result. Results are shown in Table 2 (c).
@@
Together with the results in §4.2, we observe that the performance of our framework is the following source: CellPose2D often shrinks the cell mask smaller during 2D segmentation, especially when the cell boundaries in the raw images are blurred by a light band. This shrinkage results in the loss of geometric information, as the EMD between the two masks is also affected when one mask shrinks. Moreover, the final result is also influenced by the resolution of the image stack and the average size of the cell masks. Higher resolutions amplify geometric details, enabling Geo-Wasserstein divergence to capture more nuanced information.
4.4 Ablation Analysis
| Metric | Leaf | Sepal | Pedicel | |||
| w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | |
| # Correct | 52 | 42 | 126 | 70 | 78 | 71 |
| # Unsure | 64 | 38 | 312 | 113 | 477 | 134 |
| # Incorrect | 36 | 3 | 131 | 13 | 128 | 15 |
| Accuracy | 0.5909 | 0.9333 | 0.4903 | 0.8434 | 0.3786 | 0.8256 |
| Metric | Anther | Filament | Valve | |||
| w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | |
| # Correct | 113 | 93 | 139 | 92 | 91 | 75 |
| # Unsure | 371 | 141 | 338 | 124 | 148 | 82 |
| # Incorrect | 102 | 13 | 99 | 28 | 40 | 4 |
| Accuracy | 0.5256 | 0.8807 | 0.5840 | 0.7667 | 0.6945 | 0.9494 |
Geo-Wasserstein Divergence.
@@ 4.4 Ablation Analysis
We train the classifier by discarding the Geo-Wasserstein features and evaluate on CellStitch’s flawed 3D segmentation results for labeled plant cells. Case-by-case verification results are reported in Table 3. Incorporating EMD significantly improved correction accuracy by capturing geometric differences between 2D masks that area-based metrics often overlook. While 3D topological features, such as layer-to-layer overlaps, offer limited information, geometry-aware EMD reliably detects subtle layer-to-layer 2D mask shape variations critical for distinguishing true oversegmentations from natural gaps. For instance, two geometries may have identical areas but entirely different shapes, then Geo-Wasserstein could capture this shape divergence, while the other area-based metrics failed to do so.
Statistical Features.
Following §3.2, for each candidate gap layer we summarize the layer-to-layer EMD trajectories on both sides of the gap. For each side we compute of the per-layer EMDs. The range encodes largest shift, while captures the internal trend, enabling us to assess whether the two fragments exhibit compatible statistics consistent with a single connected cell despite varying cell heights.
We evaluate robustness and sensitivity with three variants: (i) Incomplete: use only to test whether range features are necessary; (ii) Extra: augment with deciles plus and to test whether denser quantiles better capture trends; (iii) Perturbed: replace with the five quantiles , keeping dimensionality fixed to probe sensitivity to the choice of quantile levels. Results are shown in Table 4:
| Leaf | Sepal | Valve | Pedicel | Anther | Filament | |
| Statistical Feature Variants | ||||||
| mAP — Incomplete | ||||||
| mAP — Extra | ||||||
| mAP — Perturbed | ||||||
| Jac — Incomplete | ||||||
| Jac — Extra | ||||||
| Jac — Perturbed | ||||||
We find that incomplete range information (i) markedly degrades performance, whereas adding extra features (ii) yields no significant improvement over the original setup and can introduce noise. The original configuration is also robust to perturbations (iii). The number of oversegmentation candidates is comparable for (ii), (iii), and the original setup, while (i) identifies noticeably more cell-pair candidates due to less restrictions from the removed features.
@@ 5 Conclusion
We present a geometric framework for 3D cell oversegmentation, an important yet unresolved challenge that degrades the quality of 3D segmentation results. Beyond its impact on 3D cell segmentation, the proposed Geo-Wasserstein divergence is broadly applicable to segmentation problems, especially video segmentation, where temporal consistency mirrors the slice-to-slice geometric consistency exploited in 3D cell stacks: in both settings, the key question is whether regions in neighboring frames (time) or slices (depth) correspond to the same object. We hope this framework facilitates adoption by biological end-users and inspires the computer vision and geometry processing community to build upon it.
Due to space constraints, we provide additional visualizations across datasets in Appendix C of the supplementary material. Further methodological and dataset details appear in Appendix B and Appendix D, respectively.
@@ Acknowledgement
PC and YL would also like to express our appreciation to Andrew Blumberg, Abigail Hickok, and Yinuo Jin for their valuable feedback and discussions.